OpenStudy (anonymous):

WHAT ARE THE FACTOR OF 125-y raise to 3

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5^3 - y^3}\)

OpenStudy (lgbasallote):

\[\huge 125 - y^3 \leftarrow \text{difference of two cubes}\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow a^3 - b^3 = ( a -b)( a^2 + ab + b^2)}\) Here a = 5 and b = y

OpenStudy (lgbasallote):

125 = 5^3 y^3 = y^3 difference of two cubes is a^3 - b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (anonymous):

paano nio ngagawa ung example ng raise to 3

Parth (parthkohli):

Ohh both Filipino huh?

OpenStudy (anonymous):

yes are you ?

Parth (parthkohli):

No - lgbasallote is Filipino

OpenStudy (anonymous):

you ?

OpenStudy (anonymous):